OpenStudy (anonymous):

Given the function g(x) = 8x − 2, find g(−2).

OpenStudy (binary3i):

8*(-2) - 2 = -16 -2 = -18

OpenStudy (anonymous):

just put a -2 wherever you see an x and do the indicated operations... :)

OpenStudy (anonymous):

thanks you both